# A longitudinal Bayesian framework for estimating causal dose-response relationships

### v2

@@ 2.3 The generalized Bayesian bootstrap
If we consider models without the moment constraint in (3), then the solution will be the empirical distribution, i.e., $p_{i}=1/n$. Solving the moment constraint with $p_{i}=1/n$ yields a point estimate for $\xi$. If we place a prior on $p_{i}$ and allow $p_{i}$ to vary, it implicitly gives a functional form of $\xi(F)$. This is reminiscent of the classical formulation of the Bayesian bootstrap, which assumes that the data points are realizations from a multinomial model on the finite set $\left\{z_{1},\ldots,z_{n}\right\}$ with unknown probability $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$, and assuming a priori that $\mathbf{p}\sim\text{Dirichlet}\left(\alpha,\ldots,\alpha\right)$, then, a posteriori $\mathbf{p}\sim\text{Dirichlet}\left(\alpha+1,\ldots,\alpha+1\right)$. The standard Bayesian bootstrap is obtained under the improper specification $\alpha=0$ (Rubin, 1981). Therefore, if we repeatedly sample $\mathbf{p}\sim\text{Dirichlet}\left(1,\ldots,1\right)$, then
$\sum_{j=1}^{n}p_{j}\mathbf{U}^{(1)}_{j}\left(x_{j},d_{j},\gamma\right)$ $=\mathbf{0}$ (4)
@@
where the GEE, $\mathbf{U}^{(1)}(\cdot)$, is the estimating equation related to the propensity score model, and $\mathbf{U}^{(2)}(\cdot)$ is the GEE for the OR model, and $\widehat{\gamma}$ is the solution to (4). Therefore, we can obtain the posterior distribution of $\xi$ under this limiting specification (Newton and Raftery, 1994; Chamberlain and Imbens, 2003).
Motivated by this limiting case, we extend the Bayesian bootstrap to a more general Dirichlet process (DP) specification. If the approximating model, $p$, is parameterized by the true but unknown distribution, we can specify a priori $P\sim DP\left(\alpha,G_{0}\right)$ where $\alpha>0$ is the concentration parameter and $G_{0}$ is the base measure. In light of data $\left(z_{1},\ldots,z_{n}\right)$, the resulting posterior distribution of $P$ is $DP\left(\alpha_{n},G_{n}\right)$, where $\alpha_{n}=\alpha+n$ and $G_{n}(\ldotp)=\alpha G_{0}(\ldotp)\big/{(\alpha+n)}+{\sum_{j=1}^{n}\delta_{z_{j}}\left(\cdot\right)}\big/{(\alpha+n)}$. Specifically, we can generate the collection of data from the DP via $\left\{z_{j}\right\}_{j=1}^{\infty}\sim G_{n}$ and $\left\{p_{k}\right\}_{j=1}^{\infty}\sim StickBreaking(\alpha_{n})$; the standard stickbreaking algorithm (Sethuraman, 1994) generates the weights by a transformation of the collection $\{V_{k}\}_{k=1}^{\infty}$ where $V_{j}\sim Beta(1,\alpha)$ are independent, with $p_{1}=V_{1}$ and for $j=2,3,\ldots$, $p_{j}=V_{j}\prod_{k=1}^{j-1}(1-V_{k})$. Therefore, to obtain a posterior variate, we can sample $\left\{z_{j}^{s}\right\}_{j=1}^{\infty}\sim G_{n}$ and $\left\{p_{j}\right\}_{j=1}^{\infty}\sim StickBreaking(\alpha_{n})$, then solving
$\sum_{j=1}^{\infty}p_{j}\mathbf{U}^{(1)}_{j}\left(x^{s}_{j},d^{s}_{j},\gamma\right)=\mathbf{0},\sum_{j=1}^{\infty}p_{j}\mathbf{U}^{(2)}_{j}\left(z^{s}_{j},\xi,e(x_{j},\widehat{\gamma})\right)=\mathbf{0}.$ (5)
@@
  • Estimate the PS based the GEE related to $\mathbf{U}^{(1)}$ in (5) with the data $\{z^{s}_{k}\}$ and the weight $\{p_{k}^{s}\}$;
  • Solve the GEE related to $\mathbf{U}^{(2)}$ in (5) and obtain $\xi^{s}$ with the data $\{z^{s}_{k}\}$, the weight $\{p_{k}^{s}\}$ and the estimated PS;
@@ 3 Simulations
In this section, we examine the performance of the methods described in Section 2 in different aspects. We consider the following models for our simulation studies.
  • WOR-BB: The weighted estimating function approach using the GPS as the weight combined with the Bayesian bootstrap framework.
  • COV-BB: The estimation function generated by including GPS in the outcome model combined with the Bayesian bootstrap framework.
  • WOR-DP: The weighted estimating function approach using the generalized propensity score as the weight combined with the DP framework ($\alpha=5$).
  • COV-DP: The estimation function generated by including propensity scores in the outcome model combined with the DP framework ($\alpha=5$).
@@ 3.1 Example 1
In this example, we conduct a simulation study on a sample size of 100 individuals, each with 10 time points and a total of 1000 observations. We generate the data via the following relationships:
$X_{1it}\sim\mathcal{N}\left(0.2,0.1\right),X_{2it}$ $\sim\mathcal{N}\left(1,0.6\right),U_{i}\sim\mathcal{N}\left(0.2,0.1\right),D_{it}|X_{1it},X_{2i},U_{i}\sim\mathcal{N}(1+4X_{1it}+2X_{2i}+U_{i},1)$
@@
$D$ is a continuous treatment variable, and $X_{1}$, $X_{2}$ are time-varying confounders and $U$ is a time-invariant confounder for $D$ with a nonlinear relationship in $Y$. The goal of this causal setting is to estimate the APO, i.e., the outcome at a specific dose level. We consider estimating the dose-response relationship based on a GEE model for GPS with the inclusion of covariates $X_{1it}$, $X_{2it}$ only. For the COV method, we fit a GEE model for the outcome of $\log Y_{i}$ with the inclusion of covariates $D_{it}$, the three-order spline of estimated GPS. As for WOR method, we fit the outcome mean with the three-order spline of $D_{1it}$ with the GPS as the weight. In the DP method, the new data are generated from the old dataset or new data (unseen) from the base measure, $G_{0}$. We generate these new data through residuals arising from the outcome model. In this case, we first draw each pair of $\left\{y^{s^{\prime}}_{it},({x}^{s}_{it},d^{s}_{it})\right\}$, $t=1,\ldots,T_{i},i=1\ldots,N$, from the empirical distribution as the DP with $\alpha=0$, and then obtain the GPS, $\widehat{e}\left(\bar{x}^{s}_{ik}\right)$ from a mixed effect model with the inclusion of covariates $x^{s}_{1it}$, $x^{s}_{2it}$ to account for the effect of the unmeasured confounder (Graham et al., 2014), refitted to the newly sampled $\left\{({x}^{s}_{i},d^{s}_{i})\right\}$ data set. Then we simulate $y_{i}^{s}=\left\{y_{it}^{s}\right\}_{t=1}^{T}$ from a DP model with the conditional base measure $G_{0}\equiv\mathcal{N}\left(\mu({x}^{s}_{i},\widehat{e}\left(\bar{x}^{s}_{i}\right)),1\right)$, where $\mu({x}^{s}_{i},\widehat{e}\left(\bar{x}^{s}_{i}\right))$ is the fitted value from a GEE model with $\log y^{s^{\prime}}_{i}$ with the inclusion of covariates $d^{s}_{i}$, the three-order spline of estimated GPS, $\widehat{e}\left(\bar{x}^{s}_{i}\right)$. In the stick breaking construction, we truncate the sample size at $J=500$. The weight is less than $10^{-4}$ after 500 iterations and can be deemed as negligible.
We first simulate one dataset and generate 1000 posterior samples. Figure 2 shows the 1000 posterior dose-response curves with observed data points. The weighting method results in a larger variance because treatment density measurement can lead to extreme density weights and hence higher variability in the results. This issue becomes particularly obvious in the WOR approach when there are few observations in the tail range, as limited data can result in unstable weight estimates and inflated variance. In comparison, the COV approach is more stable, maintaining a comparatively lower degree of variation by exercising efficient control over the impact of fewer observations. This suggests that COV provides a more reliable estimation framework, especially in scenarios with limited sample sizes. We do not observe any noticeable difference between the BB and DP methods. Subsequently, we repeat this simulation on 1000 datasets, each consisting of 1000 entries, and evaluate the performance on three given doses. The results are shown in Table 1, with averages of the posterior means (Av Est), variances (Av Est Var) and coverage rates based on with the 2.5% and 97.5% posterior sample quantiles. The APO estimates are calculated for doses 3, 4, and 5. The APO estimates in all models are represented via a logarithmic transformation, and the results are also presented in the logarithmic scale. All Bayesian dose-response estimators have approximately unbiased estimates and desired coverage rates, despite these models involving approximations to the true nonlinear dose-response function only including the GPS and the treatment variable. When comparing the variance between the COV and WOR approaches, the WOR method exhibits greater average variability. The DP approach shows slightly lower variance in the WOR method compared to BB, and this is due to its use of a larger resampled dataset, leading to a better representation of sparsely observation regions in the original data. The BB-WOR method yields coverage around 80% which is not optimal, whereas the DP-based approach achieves coverage much closer to the nominal level. Therefore, in this case, the DP approach is preferable for the weighting-based estimator. In summary, this example suggests that the proposed double generalized Bayesian bootstrap method is suitable for handling longitudinal data with repeated measures, yielding desirable and valid uncertainty quantification.
Figure 2: Example 1: Posterior dose-response curves using WOR and COV approaches. The blue curve represents the estimated dose-response curve for one replicate. The red dots represent the observed data.
Table 1: Simulation results for the dose-response function at three given doses: average estimate, average estimated variance, coverage (%). Summary of 1000 simulation runs.
@@
Method dose = 3 dose = 4 dose = 5
@@
COV-$B$ Av Est 6.016 7.046 8.076
Av Est Var 0.004 0.003 0.004
Coverage 91.8 95.6 93.1
WOR-$B$ Av Est 6.020 7.034 8.052
Av Est Var 0.032 0.023 0.041
Coverage 84.7 84.6 86.6
COV-DP Av Est 6.020 7.049 8.079
Av Est Var 0.004 0.003 0.004
Coverage 97.7 98.5 98.3
WOR-DP Av Est 6.011 7.050 8.089
Av Est Var 0.021 0.013 0.022
Coverage 97.8 98.4 97.9
Example 2
Method dose = 3 dose = 4 dose = 5
(5.474) (6.686) (8.167)
COV Av Est 5.294 6.633 8.285
Av Est Var 0.017 0.018 0.031
Coverage 72.7 93.4 88.5
WOR Av Est 5.417 6.692 8.251
Av Est Var 0.172 0.140 0.201
Coverage 83.8 85.3 82.9
COV-DP Av Est 5.301 6.639 8.293
Av Est Var 0.017 0.018 0.029
Coverage 91.8 95.6 93.1
WOR-DP Av Est 5.459 6.723 8.271
Av Est Var 0.110 0.140 0.138
Coverage 94.8 98.8 97.5

### v3

@@ 2.3 The generalized Bayesian bootstrap
The previous section described a standard Bayesian bootstrap approach for obtaining a posterior distribution for parameters defined through estimating equations. In our setting, the model is specified through a two–stage system of GEEs, consisting of a GPS model in the first stage and an outcome model in the second stage. Specifically, let $\mathbf{U}_{1}(\cdot)$ denote the GEE associated with the GPS model and $\mathbf{U}_{2}(\cdot)$ denote the GEE for the outcome model. Under the Bayesian bootstrap, we repeatedly sample weights $\mathbf{p}=(p_{1},\ldots,p_{n})\sim\mathrm{Dirichlet}(1,\ldots,1)$, and for each draw of $\mathbf{p}$, solve the weighted estimating equations
$\sum_{j=1}^{n}p_{j}\,\mathbf{U}_{1}(x_{j},d_{j},\gamma)$ $=\mathbf{0},$ (6)
@@
where $e(x_{j};\gamma)$ denotes the GPS model, and $\widehat{\gamma}$ is the solution to (6). The resulting solution $\widehat{\xi}(\mathbf{p})$ defines a draw from the posterior distribution of $\xi$ induced by the Bayesian bootstrap. Repeating this procedure yields an empirical approximation to the posterior distribution of $\xi$. While this Bayesian bootstrap specification provides a convenient and likelihood-free way to propagate uncertainty through the two-stage GEE system, it also has an important limitation. The Bayesian bootstrap places all its prior mass on the empirical support of the observed data, implicitly assuming that the unknown data-generating distribution $F$ is discrete with atoms fixed only at the observed points. As a result, posterior inference for $\mu(d)$ is entirely driven by reweighting the observed sample and cannot extrapolate beyond the empirical support of the dose. It often tends to underestimate uncertainty in regions where data are sparse, particularly in the tails of the exposure distribution.
As proposed by Luo et al. (2023), adopting a DP prior with base measure $G_{0}$ and concentration parameter $\alpha>0$ allows posterior draws of the data-generating distribution to place mass both on the observed data and on new atoms drawn from $G_{0}$. Consequently, posterior predictive distributions for the marginal dose–response function $\mu(d)$ are no longer confined to the empirical support of the observed doses. This enables uncertainty propagation in regions of the dose space that are sparsely observed or lie near the extremes of the exposure distribution. Moreover, the base measure $G_{0}$ can carry weak structural information and can be interpreted as prior knowledge about the underlying data-generating process. In addition, the parameter $\alpha$ and $G_{0}$ can serve as a regularization mechanism, stabilizing posterior inference in regions with sparse data by shrinking toward the base measure while still allowing the data to dominate where information is abundant.
Therefore, we extend the Bayesian bootstrap to a more general DP specification. If we specify a priori $P\sim DP\left(\alpha,G_{0}\right)$ where $\alpha>0$ is the concentration parameter and $G_{0}$ is the base measure. In light of data $\left(z_{1},\ldots,z_{n}\right)$, the resulting posterior distribution of $P$ is $DP\left(\alpha_{n},G_{n}\right)$, where $\alpha_{n}=\alpha+n$ and $G_{n}(\ldotp)=\alpha G_{0}(\ldotp)\big/{(\alpha+n)}+{\sum_{j=1}^{n}\delta_{z_{j}}\left(\cdot\right)}\big/{(\alpha+n)}$. Specifically, we can generate the collection of data from the DP via $\left\{z_{j}\right\}_{j=1}^{\infty}\sim G_{n}$ and $\left\{p_{k}\right\}_{j=1}^{\infty}\sim StickBreaking(\alpha_{n})$; the standard stick-breaking algorithm (Sethuraman, 1994) generates the weights by a transformation of the collection $\{V_{k}\}_{k=1}^{\infty}$ where $V_{j}\sim Beta(1,\alpha)$ are independent, with $p_{1}=V_{1}$ and for $j=2,3,\ldots$, $p_{j}=V_{j}\prod_{k=1}^{j-1}(1-V_{k})$. Therefore, to obtain a posterior variate, we can sample $\left\{z_{j}^{s}\right\}_{j=1}^{\infty}\sim G_{n}$ and $\left\{p_{j}\right\}_{j=1}^{\infty}\sim StickBreaking(\alpha_{n})$, then solving
$\sum_{j=1}^{\infty}p_{j}\mathbf{U}_{1}\left(x^{s}_{j},d^{s}_{j},\gamma\right)=\mathbf{0},\sum_{j=1}^{\infty}p_{j}\mathbf{U}_{2}\left(z^{s}_{j},\xi,e(x_{j},\widehat{\gamma})\right)=\mathbf{0}.$ (7)
@@
In this setting, once a new dataset is generated, we first estimate the PS based the GEE related to $\mathbf{U}_{1}$, and then plug this point estimate into the GEE associated with $\mathbf{U}_{2}$. This procedure induces a random empirical measure that weights both the GPS and outcome estimating equations (with the same resampling weights) so that uncertainty is propagated coherently through the two-stage system. Both the GPS model and the outcome model are specified non-parametrically and weighted with the same stick-breaking weight. This distinguishes our approach from the previous work in Saarela et al. (2015); Liu et al. (2020). This is known as a two-step procedure (Kaplan and Chen, 2012) and is commonly used in Bayesian causal inference to avoid model feedback, as PSs should be estimated without incorporating outcome information in order to preserve their balancing properties (Rubin, 2007; McCandless et al., 2010; Saarela et al., 2015). This procedure has been shown to yield preferable estimation performance in practice (Stephens et al., 2023). They further showed that weighting both models with the same Bayesian bootstrap weights, analogous to our approach of using a common stick-breaking weight, yields valid coverage. Our Bayesian implementation in the longitudinal setting performs resampling at the subject level rather than at the observation level. Specifically, we assign a single stick-breaking weight to each study unit, which is held fixed across time. When a study unit is resampled, their entire longitudinal trajectory $\{D_{it},X_{it},Y_{it}\}_{t=1}^{T}$ is included or generated from the base measure in the resampled dataset. This subject-level resampling strategy preserves the within-subject dependence structure across repeated measurements and aligns with the marginal modeling perspective underlying GEE. By contrast, resampling study unit observations would implicitly impose independence across time within subjects and would be incompatible with the longitudinal dependence encoded in the estimating equations.
In terms of GEE-based identification under this framework, Luo et al. (2023) show that, under suitable regularity conditions, the generalized Bayesian bootstrap procedure built upon GEE or more general loss functions yield posterior distributions that concentrate asymptotically around the corresponding frequentist M-estimators. In particular, when the treatment model is correctly specified, the GEE used to define the GPS provides unbiased estimating equations for the target marginal parameters. Consequently, the resulting posterior for the APO is supported on the identified parameter space characterized by Equation (1) and concentrates around the corresponding frequentist estimand, which coincides with the true marginal dose–response function under identification assumptions stated in Section 2.1.
@@
  • Estimate the PS based the GEE related to $\mathbf{U}_{1}$ in (7) with the data $\{z^{s}_{k}\}$ and the weight $\{p_{k}^{s}\}$;
  • Solve the GEE related to $\mathbf{U}_{2}$ in (7) and obtain $\xi^{s}$ with the data $\{z^{s}_{k}\}$, the weight $\{p_{k}^{s}\}$ and the estimated PS;
@@ 3 Simulations
In this section, we examine the performance of the methods described in Section 2 in different aspects. The code for the simulation is publicly available on our GitHub page at https://github.com/yumcgill/Bayes_DoseResponse. We consider the following models for our simulation studies.
  • WOR-BB: A weighted outcome approach in which stabilized inverse GPS weights, $w_{t}=f(d)/\widehat{e}_{t}(d,X_{t})$, are incorporated into the GEE for the outcome model. The marginal density of the treatment $f(d)$ is estimated nonparametrically using a kernel density estimator. This formulation reduces variance by centering the weights around one and preventing extreme weights when the GPS is small. Uncertainty in both models is propagated by the Bayesian bootstrap framework, i.e., repeatedly resampling the weight $\mathbf{p}\sim\text{Dirichlet}\left(1,\ldots,1\right)$.
  • COV-BB: A covariate-adjusted outcome approach in which the estimated GPS, $\widehat{e}_{t}(X_{t})$, is included directly as an additional covariate in the outcome model. Posterior uncertainty is propagated using the Bayesian bootstrap framework.
  • WOR-DP: A weighted outcome approach using stabilized inverse GPS weights in the outcome GEE. Uncertainty in both models is propagated by repeatedly resampling from a stick-breaking process with concentration parameter $\alpha=5$ outlined in Section 2.3 and Algorithm 1.
  • COV-DP: A covariate-adjusted outcome approach in which the estimated GPS is included as a covariate in the outcome model. Uncertainty in both models is propagated by repeatedly resampling from a stick-breaking process with concentration parameter $\alpha=5$.
  • Bayes-MSM: A Bayesian MSM in which the causal dose–response function is modeled directly as a parametric function of the dose. Stabilized inverse GPS weights, $w_{t}=f(d)/\widehat{e}_{t}(d,X_{t})$, are incorporated through a weighted likelihood to remove time-dependent confounding. The GPS and outcome models are both modeled parametrically. The outcome model is specified as a cubic spline regression of the outcome on dose, with non-informative priors placed on all regression coefficients and variance parameters. Posterior inference for both GPS and outcome models is obtained via MCMC using Stan (Stan Development Team, 2025). Note this formulation is not a Bayesian analysis based on a fully specified joint likelihood factorization for all observed processes with independent priors on treatment and outcome model parameters (see discussion in Saarela et al., 2015); rather, it follows a Bayesian working-model approach in which weighting is used to target the causal estimand directly.
Specifically, we outline the procedure for generating posterior samples of the APO at a given dose level $d$ using the proposed DP–based method:
  1. 1.
    Resampling under the DP prior. For each Monte Carlo iteration $s=1,\ldots,S$, we first resampled with replacement from the original dataset, obtain the new dataset with the number of study units $N=100$. This step can be viewed as the DP with $\alpha=0$. For each resampled study unit, a random draw determines whether the study unit’s outcome, $y$, is taken directly from the base distribution, with probabilities $\alpha/(\alpha+n)$. For study units drawn from the base distribution, new outcome values are simulated from a normal distribution with mean given by the model-based predicted value and a standard deviation of 0.5. For each study unit, the predicted mean is derived from the outcome model specified in the corresponding simulation example.
  2. 2.
    Weight generation via stick-breaking. Given the resampled data, a corresponding set of weights $\{p_{k}^{(s)}\}$ is drawn from a stick-breaking process with concentration parameter $\alpha_{n}=\alpha+n$. These weights are assigned to the resampled study units and carried forward into the analysis.
  3. 3.
    GPS estimation (first stage). Using the resampled data from Step 1 and weights from Step 2, the GPS is estimated by solving GEE associated with $\mathbf{U}_{1}$. This stage is performed independently of the outcome model in order to avoid feedback from the outcome into the PS estimation.
  4. 4.
    Outcome model estimation (second stage). Conditional on the estimated GPS, the GEE associated with $\mathbf{U}_{2}$ in Equation (2) is solved using the same resampled data and weights, yielding a posterior draw $\xi^{(s)}$ of the structural parameters.
  5. 5.
    Causal estimand computation. Based on the parameter draw $\xi^{(s)}$ and the estimated GPS, we can predict outcome, $\zeta^{(s)}(d)$, at specific dose level $d$. Repeating the above steps for $s=1,\ldots,S$ yields posterior samples
    $\{\zeta^{(1)}(d),\ldots,\zeta^{(S)}(d)\},$
    which are used to summarize uncertainty in the APO through posterior means, credible intervals, and coverage rates.
@@ 3.1 Example 1
In this example, we conduct a simulation study on a sample size of 100 study units, each with 10 time points and a total of 1000 observations. We generate the data via the following relationships:
$X_{1it}\sim\mathcal{N}\left(0.2,0.1\right),X_{2it}$ $\sim\mathcal{N}\left(1,0.6\right),U_{i}\sim\mathcal{N}\left(0.2,0.1\right)$
$D_{it}|X_{1it},X_{2i},U_{i}$ $\sim\mathcal{N}(5+4X_{1it}+2X_{2i}+U_{i},1)$
@@
$D$ is a continuous treatment variable, and $X_{1}$, $X_{2}$ are time-varying confounders and $U$ is a time-invariant confounder for $D$ with a nonlinear relationship in $Y$. The goal of this causal setting is to estimate the APO, i.e., the outcome at a specific dose level. For the GEE associated with the first stage of the GPS estimation, $\mathbf{U}_{1}$, we include only the covariates $X_{1it}$ and $X_{2it}$. For the GEE associated with the outcome model, $\mathbf{U}_{2}$, we use the outcome in log scale, $\log Y_{i}$. Specifically, in the covariate-adjusted (COV) method, we include $D_{it}$ and a cubic spline of the estimated GPS, allowing for a flexible, nonlinear relationship between the GPS and the outcome. In the weighted outcome regression (WOR) method, we include a cubic spline of $D_{it}$ only. For the DP-based method, the predicted mean is obtained from a GEE outcome model that includes $D_{it}$ and a cubic spline of the estimated GPS.
We first simulate one dataset and generate 1000 posterior samples. Figure 2 shows the 1000 posterior dose-response curves with observed data points. All methods recover the general monotonic increasing relationship between dose and outcome, but there are differences in tail behavior and uncertainty quantification. The WOR-based approaches exhibit some deviation from the true dose–response curve, particularly in the high dose region. However, these regions are characterized by larger variance, leading to wider credible intervals that maintain coverage of the true curve. The DP-WOR method yields slightly wider intervals than BB-WOR, reflecting the additional uncertainty introduced by the more flexible prior on the data-generating distribution. In contrast, the COV-based approaches produce smoother dose–response estimates with relatively narrow credible intervals and remain close to the true curve even in regions with limited data support. This suggests that the COV-based formulation provides a more stable and reliable estimation framework, particularly in settings with small sample sizes. Finally, the Bayes-MSM approach also yields smooth estimates with comparatively tight credible intervals, but shows noticeable departures from the true curve in both tails, highlighting its sensitivity to parametric assumptions when extrapolating beyond well-supported regions of the data.
Subsequently, we repeat this simulation on 1000 datasets, each consisting of 1000 entries, and evaluate the performance on three given doses. The results are shown in Table 1, with averages of the posterior means (Est), standard deviations (SD) and coverage rates based on with the 2.5% and 97.5% posterior sample quantiles. The true APO at dose $d$ was estimated by Monte Carlo integration with 100,000 observations, corresponding to 10,000 study units with 10 measurements each. The APO estimates are calculated for doses 6, 8, and 10. The APO estimates in all models are represented via a logarithmic transformation, and the results are also presented in the logarithmic scale. All methods yield the average of the posterior means close to the true values. The COV-based approaches are more efficient than the WOR-based ones, with smaller standard deviations. The DP-based methods attain near-nominal coverage across all doses considered, whereas COV-BB achieves nominal coverage only at doses with adequate data support and exhibits undercoverage at sparsely observed doses. The WOR-BB method shows consistently lower coverage across doses, while the Bayes-MSM displays undercoverage at doses with limited data support. Overall, this example indicates that the proposed DP-based methods are well suited for longitudinal data with repeated measures, providing valid uncertainty quantification even in low-data regions.
Figure 2: Example 1: Posterior predictive dose–response curves for the Gaussian outcome estimated using the Bayesian bootstrap, Dirichlet process, and Bayesian parametric MSM approaches.
Table 1: Simulation results for the dose–response function at three dose levels, reporting the average posterior mean (Est) with associated standard deviation (SD), and empirical coverage rate (Coverage, %). Results are summarized over 1,000 simulation replicates.
Example 1 (Results are presented in the logarithmic scale)
Method dose = 6 dose = 8 dose = 10
(9.046) (11.046) (13.046)
COV-$B$ Est 8.987 11.045 13.104
SD 0.046 0.036 0.048
Coverage 75.4 94.4 75.7
WOR-$B$ Est 8.959 11.037 13.118
SD 0.217 0.145 0.205
Coverage 77.6 84.3 78.3
COV-DP Est 8.987 11.045 13.103
SD 0.086 0.055 0.088
Coverage 98.2 98.7 97.4
WOR-DP Est 8.964 11.039 13.115
SD 0.188 0.128 0.176
Coverage 96.9 98.8 96.7
Bayes-MSM Est 8.994 11.043 13.097
SD 0.091 0.068 0.096
Coverage 84.5 92.3 82.2
@@
Method dose = 6 dose = 8 dose = 10
@@
COV-$B$ Est 9.175 13.737 20.570
SD 0.128 0.131 0.195
Coverage 93.7 93.9 91.8
WOR-$B$ Est 9.097 13.739 20.686
SD 0.547 0.498 0.972
Coverage 82.8 85.5 81.8
COV-DP Est 8.879 13.240 19.817
SD 0.202 0.182 0.381
Coverage 98.2 91.5 98.0
WOR-DP Est 8.763 13.220 19.879
SD 0.469 0.454 0.788
Coverage 94.9 94.2 97.1
Bayes-MSM Est 9.103 13.686 20.494
SD 0.286 0.261 0.446
Coverage 81.2 91.1 90.4